# Exploring the Potential of Encoder-free Architectures in 3D LMMs

### v2

@@ 3.1 Overall Architecture
Token Embedding. We first remove the encoder of PointLLM and adopt the original token embedding (Yu et al., 2022). However, the coarse structural design results in a significant performance degradation, as observed in Table 1, where the GPT-4 scores for the classification and captioning tasks decrease by 17.5% and 10.48%, respectively. To mitigate excessive information loss and provide refined local features to the LLM, we adopt a small network with a limited number of parameters, which is a lightweight variant of Point-PN (Zhang et al., 2023b). Specifically, for the input $\{P_{i}\}_{i=1}^{N}$, we apply Farthest Point Sampling (FPS) for downsampling the number of points, k-Nearest Neighbors (k-NN) with group size k for local aggregation, and learnable linear layers for feature encoding. After a series of repetitive operations and the projection layer, we transform the point cloud into high-dimensional vectors $\{F_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times D_{1}}$. In Table 1, we experiment with token embedding at different depths and find that three layers yield the best performance, while two layers fail to capture complex point features and four layers introduce noise.
Table 3: LLM-embedded Semantic Encoding. In pre-training, we explore the effects of different self-supervised learning losses targeting point tokens. $\Psi$ and $\Phi$ denote mask ratios of 60% and 30%, respectively. Subscripts $patch$ and $feat$ indicate loss targets. For Hybrid Semantic Loss, the subscripts $patch$ and $feat$ refer to the masked modeling target, with reconstruction targeting the corresponding $feat$ and $patch$.
@@
Masked Modeling LossfeatΨ 50.00 45.80 46.29
Masked Modeling LossfeatΦ 49.50 47.35 47.93
Reconstruction Losspatch 49.50 46.96 47.33
Reconstruction Lossfeat 48.50 45.95 47.18
Contrastive Loss 43.50 42.91 44.77
Knowledge Distillation Loss 49.50 45.43 47.09
Hybrid Semantic Losspatch 50.50 46.84 47.59
Hybrid Semantic Lossfeat 53.00 48.85 48.00
Further 3D Encoding & Alignment. We discover that the absence of the encoder results in a lack of context modeling in point cloud feature processing. Therefore, we attempt to have the early layers of the LLM take on the encoder’s role in capturing global interactions of features, further encoding the point cloud features. In the pre-training stage, we set the first K layers of the frozen LLM to be learnable, utilizing the self-attention mechanism to capture global geometric structures. In the first K layers, 3D tokens and text tokens interact and align naturally within a shared semantic space. Meanwhile, we experiment with different learning rates. As shown in Table 1, a smaller learning rate yields better results by stabilizing early layer optimization. Based on the designed token embedding module, setting the first four layers to be learnable yields the best results, as it effectively encodes low-level features into high-level representations with considerable computational efficiency.
3.2 LLM-embedded Semantic Encoding
The lack of the 3D encoder results in insufficient encoding of point cloud semantic information, which greatly hinders the LLM to understand the structural details of point clouds. Most existing 3D encoders use self-supervised losses to embed the high-level semantics of point clouds into the transformer, primarily categorized into four types: Masked Modeling Loss (Pang et al., 2022), Reconstruction Loss (Qi et al., 2023), Contrastive Loss (Khosla et al., 2020), and Knowledge Distillation Loss (Zhang et al., 2023a). Based on the proposed token embedding module and LLM learnable early layers, we implement and evaluate the effects of these losses on the encoder-free 3D LMM in the pre-training stage, as described in Figure 3. Finally, we propose the Hybrid Semantic Loss, which assists the LLM to learn the relationship between local spatial information in the point clouds and grasp the high-level 3D semantics.
Masked Modeling Loss. In the pre-training stage, we apply the Masked Modeling Loss to the point tokens processed by the LLM, as shown in Figure 3 (a). Through the Farthest Point Sampling (FPS) and k-Nearest Neighbors (k-NN) algorithms in the token embedding, the point clouds $\{P_{i}\}_{i=1}^{N}$ are divided into point patches $\{G_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times k\times 3}$ and the corresponding point tokens $\{F_{i}\}_{i=1}^{M}$. We randomly mask the point tokens with a masking ratio $r$, and replace them with learnable tokens. The masked feature tokens can be denoted as $\{F_{\text{gt}_{i}}\}_{i=1}^{M*r}$, which serve as the ground truth for the loss computation. After the learnable tokens are concatenated with visible tokens and processed by the LLM, a linear layer is applied to predict the point tokens $\{F_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times D_{1}}$, and the Mean Squared Error (MSE) is computed between $F_{\text{pre}}$ and $F_{\text{gt}}$. The optimization can be written as
The specific process of applying Masked Modeling to point patches $G$ is detailed in Appendix B.1.
where $a=G_{\text{pre}}$, $b=G$. The procedure for reconstructing feature F is detailed in Appendix B.1.
Contrastive Loss. We conduct contrastive learning (Khosla et al., 2020) at the point cloud level, where we contrast two transformed versions of the point cloud in the Figure 3 (c). Given a sampled point cloud $\{P_{i}\}_{i=1}^{N}$, we apply two random geometric transformations $T_{1}$ and $T_{2}$, including rotation and translation, to obtain $P_{T1}$ and $P_{T2}$. The two augmented point clouds are separately paired with the original text query and processed through the LLM to obtain their respective feature tokens $F_{T1}\in\mathbb{R}^{M\times D_{1}}$ and $F_{T2}\in\mathbb{R}^{M\times D_{1}}$. Within the mini-batch, the two feature tokens derived from the same point cloud serve as positive pairs, while they are considered negative pairs with other point clouds. Using NCESoftmaxLoss, we aim to maximize the similarity of positive pairs and minimize the similarity of negative pairs, encouraging the LLM to learn geometric equivariance of point clouds. The $\mathcal{L}_{\text{contrast}}$ is:
$\frac{1}{B}\sum_{i=1}^{B}\left(-\log\frac{\exp(\mathbf{F}_{T1_{i}}\cdot\mathbf{F}_{T2_{i}}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{F}_{T1_{i}}\cdot\mathbf{F}_{T2_{j}}/\tau)}\right),$ (3)
where B stands for the training batch size.
Knowledge Distillation Loss. We select the powerful Uni3D-L (Zhou et al., 2023) as the teacher encoder, input the point cloud into the 3D encoder, and obtain the output feature $F_{\text{teacher}}\in\mathbb{R}^{M\times D_{2}}$. The Mean Squared Error (MSE) between the LLM output tokens $F_{\text{student}}$ and $F_{\text{teacher}}$ is computed to align $F_{\text{student}}$ as closely as possible to $F_{\text{teacher}}$, thereby transferring the knowledge embedded in the 3D encoder to the LLM. By obtaining additional supervision from the Uni3D, the LLM better captures the complex structures in the point cloud data, as displayed in Figure 3 (d). The objective function is defined as
@@ 3.2 LLM-embedded Semantic Encoding
$l$=1 52.00 48.74 48.04
$l$=2 50.80 46.60 47.95
$l$=3 52.60 48.92 47.99
$l$=4 49.10 45.51 46.81
$H$=2 53.58 49.20 48.43
$H$=4 52.40 48.41 47.75
$H$=8 52.06 48.95 47.90
+ Self-Attn. 55.10 50.98 48.65
Experiments and Insights. As shown in Table 3, we compare the effects of common self-supervised learning losses in the pre-training stage, where they are summed with the LLM cross-entropy loss (Touvron et al., 2023), each with a coefficient of 1. The observations are summarized as below:
  • The point cloud self-supervised losses generally benefit the encoder-free 3D LMM. Compared to previous experimental results, where the GPT scores for the classification and captioning tasks are 49.11% and 45.39%, the self-supervised losses bring about the significant improvements. This is because the self-supervised learning loss forces transformations on the complex point clouds through certain task design. This encourages the LLM to not simply memorize specific point cloud data but to learn the underlying geometric relationships and high-level semantic information.
  • Among the self-supervised learning losses, the Masked Modeling Loss demonstrates the strongest performance improvement. It achieves GPT-4 scores of 49.5% and 47.35% for classification and captioning tasks, respectively. The application of the masked modeling to the point features facilitates the embedding of high-level semantics from point clouds into the LLM. However, a higher mask ratio increases training difficulty, with 60% performing worse than 30%. In addition, explicitly reconstructing point patches helps capture complex structures and critical details in point clouds. Knowledge Distillation Loss falls short compared to the first two losses. Finally, Contrastive Loss, which fails to extract the detailed semantics, achieves the lowest performance.
Hybrid Semantic Loss. Based on the experimental results above, we propose the self-supervised learning loss specifically designed for the encoder-free 3D LMM—Hybrid Semantic Loss, as showcased in Figure 3 (e). We apply a masking ratio $r$ to randomly mask point tokens from the token embedding. The masked tokens and the corresponding patches are referred to as $\{F_{\text{mask}_{i}}\}_{i=1}^{M*r}$ and $\{G_{\text{mask}_{i}}\}_{i=1}^{M*r}$, respectively. The remaining tokens are denoted as$\{F_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$ and $\{G_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$. Considering the autoregressive nature of the LLM and the unordered attribute of point clouds, we directly concatenate learnable tokens $\{F_{\text{learn}_{i}}\}_{i=1}^{M*r}$ to the end of $F_{\text{vis}}$, replacing the masked tokens. For the masked portion, we adopt masked modeling, and for the visible portion, we use the reconstruction strategy. After passing point tokens through the LLM, we compute the MSE between $F_{\text{learn}}$ and $F_{\text{mask}}$. The visible features $F_{\text{vis}}$ are transformed into $G_{\text{pred}}$, and the $L_{2}$ Chamfer distance is computed between $G_{\text{pred}}$ and $G_{\text{vis}}$. These two are added to the original cross-entropy loss with coefficients all equal to 1. This approach not only embeds high-level semantics into the LLM but also ensures geometric consistency throughout the point cloud learning process. With a 30% mask ratio and per-layer positional encoding of point tokens, it achieves 53.00% and 48.85% on the classification and captioning tasks, respectively. The inverse modeling process is described in Appendix B.1.
@@
Table 5: Comparison of different models on various 3D understanding tasks. A primary focus is placed on GPT-4 evaluation, along with data-driven metrics (Sentence-BERT and SimCSE).
@@
InstructBLIP-7B(Dai et al., 2023) 45.34 47.41 48.48 4.27 8.28 12.99 43.50
InstructBLIP-13B(Dai et al., 2023) 44.97 45.90 48.86 4.65 8.85 13.23 34.25
LLaVA-7B(Liu et al., 2024) 46.71 45.61 47.10 3.64 7.70 12.14 50.00
LLaVA-13B(Liu et al., 2024) 38.28 46.37 45.90 4.02 8.15 12.58 51.75 47.90
3D-LLM(Hong et al., 2023) 33.42 44.48 43.68 16.91 19.48 19.73 45.25
PointLLM-7B(Xu et al., 2023) 44.85 47.47 48.55 3.87 7.30 11.92 53.00 41.20
PointLLM-13B(Xu et al., 2023) 48.15 47.91 49.12 3.83 7.23 12.26 54.00 46.60
ShapeLLM-7B(Qi et al., 2024) 46.92 48.20 49.23 54.50 47.40
ShapeLLM-13B(Qi et al., 2024) 48.94 48.52 49.98 54.00 53.10
Enel-7B 50.98 48.65 49.36 3.85 7.20 12.60 55.10 43.10
Enel-13B 52.87 48.85 50.03 3.72 7.25 12.31 55.30 48.20
3.3 Hierarchical Geometry Aggregation
3D encoders are designed with specific structures tailored for point clouds, such as local-to-global hierarchy (Zhang et al., 2022) for exploring the geometric structure of the point cloud. However, in encoder-free architectures, the LLM itself does not have an explicit local modeling module. The self-attention mechanism is intended for modeling global interactions. Therefore, building upon the proposed Hybrid Semantic Loss, we explore in the instruction tuning stage how to enable the LLM to actively perceive 3D local details and complement the learned global semantics. To this end, we propose the Hierarchical Geometry Aggregation strategy.
Implementation Details. As depicted in Figure 4, from the LLM second layer, the input point tokens $\{F_{\text{input}_{i}}\}_{i=1}^{M}$, based on their corresponding coordinates $\{P_{\text{input}_{i}}\}_{i=1}^{M}$, are downsampled using the Farthest Point Sampling (FPS), reducing the token number from $M$ to $M/2$, denoted as $F_{\text{input}}^{c}$, which serve as the local centers. Then, using the k-Nearest Neighbor (k-NN) algorithm, we obtain the neighboring points $F_{\text{input}}^{n}\in\mathbb{R}^{M/2\times k\times D_{1}}$ for the center points. For $F_{\text{input}}^{n}$, we employ the gated self-attention mechanism for intra-group interactions, grasping the local geometric structure. We multiply the self-attention output by a learnable parameter initialized from zero to adaptively adjust the required knowledge. We formulate it as
On top of this, we apply pooling to fuse the features ${F_{\text{input}}^{n}}^{\prime}$ within each neighbor and add them to the original center tokens $F_{\text{input}}^{c}$, yielding aggregated tokens $\{F_{\text{agg}_{i}}^{1}\}_{i=1}^{M/2}$, formulated as
$\begin{split}{F_{\text{agg}}^{1}}=\text{Pooling}({F_{\text{input}}^{n}}^{\prime})+{F_{\text{input}}^{c}}.\end{split}$ (6)
Then we perform $l-1$ iterations of geometry aggregation, resulting in $\{F_{\text{agg}_{i}}^{l}\}_{i=1}^{M/2^{l}}$. To ensure that the LLM fully extracts the local information, we choose to perform further semantic modeling using $H$ LLM layers after aggregation operations. This allows the model to learn the interactions between local information while preventing the loss of fine-grained geometric details. Subsequently, from the $L$th layer, we perform $l$ iterations of geometry propagation. Following the approach of PointNet++ (Qi et al., 2017), we propagate the aggregated features $F_{\text{agg}}^{l}$ from the local center points to their surrounding $k$ neighboring points, generating $\{F_{\text{pro}_{i}}^{1}\}_{i=1}^{M/2^{(l-1)}}$. After $l$ iterations, we obtain point tokens of length $M$, which are then processed by the remaining LLM layers.
@@ 3.3 Hierarchical Geometry Aggregation
  • The best performance is achieved when $l$ is set to 3. As observed in Table 4, performing triple aggregation and propagation operations achieves 48.92% and 52.60% performance on the captioning and classification tasks, respectively. Fewer aggregation layers $l$ limit local geometric capture, while too many overly simplify spatial relationships. Setting $l=3$ achieves a balanced modeling of local and global structures.
  • Compared to setting $H$ to 4 or 8, the highest performance is achieved when $H$ is set to 2. It reaches 53.58% and 49.20% on the classification and captioning tasks, respectively. The excessive number of LLM layers between aggregation and propagation can lead to the oversmoothing of the aggregated local information, resulting in the loss of local structural details.
  • The gated self-attention mechanism effectively improves performance, reaching 55.10% and 50.98% on classification and captioning tasks, respectively. The adaptive control of attention output ensures that global contextual information is utilized only when necessary, preventing it from disrupting local geometric structures. Additionally, it allows the model to adjust to different tasks.

### v4

@@ 3.2 LLM-embedded Semantic Encoding
Contrastive Loss. We conduct contrastive learning Khosla et al. (2020) at the point cloud level, where we contrast two transformed versions of the point cloud in the Figure 3 (c). Given a sampled point cloud $\{P_{i}\}_{i=1}^{N}$, we apply two random geometric transformations $T_{1}$ and $T_{2}$, including rotation and translation, to obtain $P_{T1}$ and $P_{T2}$. The two augmented point clouds are separately paired with the original text query and processed through the LLM to obtain their respective feature tokens $F_{T1}\in\mathbb{R}^{M\times D_{1}}$ and $F_{T2}\in\mathbb{R}^{M\times D_{1}}$. Within the mini-batch, the two feature tokens derived from the same point cloud serve as positive pairs, while they are considered negative pairs with other point clouds. Using NCESoftmaxLoss, we aim to maximize the similarity of positive pairs and minimize the similarity of negative pairs, encouraging the LLM to learn geometric equivariance of point clouds. The $\mathcal{L}_{\text{contrast}}$ is shown as below, where B stands for the training batch size.
$\frac{1}{B}\sum_{i=1}^{B}\left(-\log\frac{\exp(\mathbf{F}_{T1_{i}}\cdot\mathbf{F}_{T2_{i}}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{F}_{T1_{i}}\cdot\mathbf{F}_{T2_{j}}/\tau)}\right).$ (3)
Knowledge Distillation Loss. We select the powerful Uni3D-L Zhou et al. (2023) as the teacher encoder, input the point cloud into the 3D encoder, and obtain the output feature $F_{\text{teacher}}\in\mathbb{R}^{M\times D_{2}}$. The Mean Squared Error (MSE) between the LLM output tokens $F_{\text{student}}$ and $F_{\text{teacher}}$ is computed to align $F_{\text{student}}$ as closely as possible to $F_{\text{teacher}}$, thereby transferring the knowledge embedded in the 3D encoder to the LLM. By obtaining additional supervision from the Uni3D, the LLM better captures the complex structures in the point cloud data, as displayed in Figure 3 (d). The objective function is:
Table 3: LLM-embedded Semantic Encoding. In pre-training, we explore the effects of different self-supervised learning losses targeting point tokens. $\Psi$ and $\Phi$ denote mask ratios of 60% and 30%, respectively. Subscripts patchpatch and featfeat indicate loss targets. For Hybrid Semantic Loss, the subscripts patchpatch and featfeat refer to the masked modeling target, with reconstruction targeting the corresponding featfeat and patchpatch.
@@
Masked Modeling LossfeatΨ 48.50 43.90 45.30
Masked Modeling LossfeatΦ 48.50 45.85 46.93
Reconstruction Losspatch 48.00 45.56 46.33
Reconstruction Lossfeat 47.50 44.05 46.18
Contrastive Loss 42.50 41.21 43.77
Knowledge Distillation Loss 48.00 43.87 46.09
Hybrid Semantic Losspatch 50.00 45.24 46.59
Hybrid Semantic Lossfeat 52.00 47.65 47.30
Experiments and Insights. As shown in Table 3, we compare the effects of common self-supervised learning losses in the pre-training stage, where they are summed with the LLM cross-entropy loss Touvron et al. (2023), each with a coefficient of 1. The observations are summarized as below:
  • The point cloud self-supervised losses generally benefit the encoder-free 3D LMM. Compared to previous experimental results, where the GPT scores for the classification and captioning tasks are 47.90% and 43.50%, the self-supervised losses bring about the significant improvements. This is because the self-supervised learning loss forces transformations on the complex point clouds through certain task design. This encourages the LLM to not simply memorize specific point cloud data but to learn the underlying geometric relationships and high-level semantic information.
  • Among the self-supervised learning losses, the Masked Modeling Loss demonstrates the strongest performance improvement. It achieves GPT-4 scores of 48.5% and 45.85% for classification and captioning tasks, respectively. The application of the masked modeling to the point features facilitates the embedding of high-level semantics from point clouds into the LLM. However, a higher mask ratio increases training difficulty, with 60% performing worse than 30%. In addition, explicitly reconstructing point patches helps capture complex structures and critical details in point clouds. Knowledge Distillation Loss falls short compared to the first two losses. Finally, Contrastive Loss, which fails to extract the detailed semantics, achieves the lowest performance.
[…]
Our motivation arises from the observation that complex objectives, such as KD and contrastive learning, impose significant computational overhead yet often yield marginal gains compared to intrinsic data modeling losses like masked modeling. To address this, we propose the Hybrid Semantic Loss, which resolves the structural mismatch between 3D data and LLMs by exploiting two key properties: (1) the permutation invariance of point clouds, allowing learnable tokens to be appended after visible tokens without positional restoration; and (2) the encoder-free architecture, where 3D tokens are integrated into a causally-masked LLM instead of a bidirectionally-masked 3D encoder, fundamentally altering information flow between visible and masked tokens, enabling visible tokens to learn harder objectives while learnable tokens focus on lightweight reconstruction.
3.3 Hierarchical Geometry Aggregation
3D encoders are designed with specific structures tailored for point clouds, such as local-to-global hierarchy Zhang et al. (2022) for exploring the geometric structure of the point cloud. However, in encoder-free architectures, the LLM itself does not have an explicit local modeling module. The self-attention mechanism is intended for modeling global interactions. Therefore, building upon the proposed Hybrid Semantic Loss, we explore in the instruction tuning stage how to enable the LLM to actively perceive 3D local details and complement the learned global semantics. To this end, we propose the Hierarchical Geometry Aggregation strategy.
@@ 3.3 Hierarchical Geometry Aggregation
ll=1 52.50 48.70 48.07
ll=2 51.00 46.67 48.12
ll=3 53.00 48.93 48.06
ll=4 45.00 45.48 46.90
HH=2 54.25 49.56 48.52
HH=4 52.50 48.61 47.81
HH=8 52.25 48.95 47.90
+ Self-Attn. 55.55 51.03 48.79
Implementation Details. As depicted in Figure 4, from the LLM second layer, the input point tokens $\{F_{\text{input}_{i}}\}_{i=1}^{M}$, based on their corresponding coordinates $\{P_{\text{input}_{i}}\}_{i=1}^{M}$, are grouped by Dynamic Grid Sampling. The grid size follows a cumulative scaling strategy across aggregation layers. At the $i$-th aggregation layer, the grid size is:
$s_{i}=\alpha\cdot e^{\sum_{j=1}^{i}\beta_{j}},\qquad\beta_{j}=\gamma\cdot\tanh(\theta_{j})+\beta_{\text{ctr}},$ (5)
where $\alpha=0.02$ m and $s_{i}\in[s_{\min},s_{\max}]=[0.02,1]$ m. To ensure the cumulative scaling stays within bounds across $l$ aggregation layers, we set:
$\gamma=\frac{\ln\left(\frac{s_{\max}}{\alpha}\right)-\ln\left(\frac{s_{\min}}{\alpha}\right)}{2l},\qquad\beta_{\text{ctr}}=\frac{\ln\left(\frac{s_{\max}}{\alpha}\right)+\ln\left(\frac{s_{\min}}{\alpha}\right)}{2l},$ (6)
where $l$ is the total number of aggregation layers. Each $\theta_{j}$ is randomly initialized from a standard normal distribution. Points within the same grid cell form local neighbors, with the set of all neighbors denoted as $\mathcal{G}_{i}$ having cardinality $M_{i}$. The neighborhood features $F_{\text{input}}^{n}\in\mathbb{R}^{M_{i}\times k\times D_{1}}$ are then collected, where $k$ denotes the maximum number of points across all cells. To handle varying point numbers across grid cells, we employ a padding strategy: for cells with fewer than $k$ points, we compute the mean-pooled feature of existing points and concatenate it repeatedly until reaching $k$ points per cell. For $F_{\text{input}}^{n}$, we employ the gated self-attention mechanism for intra-group interactions, grasping the local geometric structure. We multiply the self-attention output by a learnable parameter initialized from zero to adaptively adjust the required knowledge. We formulate it as
On top of this, we apply pooling to fuse the features $F_{\text{input}}^{n}$$^{\prime}$ within each neighbor, yielding aggregated tokens $\{F_{\text{agg}_{j}}^{i}\}_{j=1}^{M_{i}}$, formulated as
$\begin{split}{F_{\text{agg}}^{i}}=\text{MeanPooling}({F_{\text{input}}^{n}}^{\prime}).\end{split}$ (8)
We perform $l$ iterations of geometry aggregation, resulting in $\{F_{\text{agg}_{i}}^{l}\}_{i=1}^{M_{l}}$. To ensure that the LLM fully extracts the local information, we choose to perform further semantic modeling using $H$ LLM layers after aggregation operations. This allows the model to learn the interactions between local information while preventing the loss of fine-grained geometric details. Subsequently, from the $L$th layer, we perform $l$ iterations of geometry propagation. Following the grid unpooling strategy, we use the point-to-grid mappings to propagate the aggregated features $F_{\mathrm{agg}}^{l}$ from each grid cell back to its corresponding set of points, generating $\{\,F^{1}_{\mathrm{pro}_{i}}\,\}_{i=1}^{M_{l-1}}$. After $l$ iterations, we obtain point tokens of length $M$, which are then processed by the remaining LLM layers. After processing through H additional LLM layers, the geometry aggregation and propagation process is repeated.
@@
Table 5: Comparison of different models on various 3D understanding tasks. A primary focus is placed on GPT-4 evaluation, along with data-driven metrics (Sentence-BERT). The * indicates the Qwen2.5-7B LLM base and the ShapeLLM training data. The $\alpha$ denotes reproduced results. denotes the model is implemented based on the ShapeLLM baseline.
@@
InstructBLIP-7BDai et al. (2023) 45.34 47.41 48.48 4.27 8.28 12.99 43.50
InstructBLIP-13BDai et al. (2023) 44.97 45.90 48.86 4.65 8.85 13.23 34.25
LLaVA-7BLiu et al. (2024) 46.71 45.61 47.10 3.64 7.70 12.14 50.00
LLaVA-13BLiu et al. (2024) 38.28 46.37 45.90 4.02 8.15 12.58 51.75 47.90
PointGPTChen et al. (2023) 11.60
Uni3DZhou et al. (2023) 47.20
3D-LLMHong et al. (2023) 33.42 44.48 43.68 16.91 19.48 19.73 45.25
PointLLM-7BXu et al. (2023b) 44.85 47.47 48.55 3.87 7.30 11.92 53.00 41.20
PointLLM-13BXu et al. (2023b) 48.15 47.91 49.12 3.83 7.23 12.26 54.00 46.60
ShapeLLM-7BQi et al. (2024) 46.92 48.20 49.23 54.50 47.40
ShapeLLM-13BQi et al. (2024) 48.94 48.52 49.98 54.00 53.10
MiniGPT-3Dα Tang et al. (2024c) 52.49 48.73 49.26 54.50 43.60
PointLLM-PiSA-7BGuo et al. (2025a) 48.63 48.47 49.08 3.80 7.25 12.38 54.50 42.90
PointLLM-PiSA-13BGuo et al. (2025a) 50.52 48.60 49.64 3.75 7.84 12.56 55.00 46.80
Enel-7B 51.03 48.79 49.52 3.91 7.20 12.68 55.55 43.80
Enel-7B 53.26 48.75 49.94 - - - 56.00 48.90
Enel-13B 53.24 48.92 50.17 3.72 7.89 12.31 56.00 48.50
Enel-13B 54.78 49.37 50.69 - - - 56.00 54.80
Enel-7B 57.91 49.90 51.84 5.32 8.58 13.98 61.00 55.20
@@
  • The best performance is achieved when $l$ is set to 3. As shown in Table 4, performing three aggregation and propagation operations achieves 48.93% and 53.00% performance on captioning and classification tasks, respectively. Fewer aggregation layers limit the capture of local geometric information, while too many layers oversimplify spatial relationships. Setting $l=3$ achieves balanced modeling of local and global structures and realizes sampling ratio of approximately 1/8.
  • Compared to setting $H$ to 4 or 8, the highest performance is achieved when $H$ is set to 2. It reaches 54.25% and 49.56% on the classification and captioning tasks, respectively. The excessive number of LLM layers between aggregation and propagation can lead to the oversmoothing of the aggregated local information, resulting in the loss of local structural details.
  • The gated self-attention mechanism effectively improves performance, reaching 55.55% and 51.03% on classification and captioning tasks, respectively. The adaptive control of attention output ensures that global contextual information is utilized only when necessary, preventing it from disrupting local geometric structures. Additionally, it allows the model to adjust to different tasks.